OpenStudy (anonymous):

To find the perimeter of a rectangle you use the formula: P = 2(l + w). l is the length of the rectangle. w is the width of the rectangle. Which equation would you use to find the value of w if you were given l and P?

OpenStudy (bradely):

P = 2(l + w). Divide each side by 2 P/2 =l+w Subtract l from each side w = P/2 - l Source: http://www.mathskey.com/question2answer/